Use either the fundamental counting principle or the permutation formulas (or both) to answer the following:
How many seven-digit phone numbers can be formed if the first digit cannot be 0 and
a.) repetition of digits is not permitted?
b.) if repetition of digits is permitted?





Answer by oscargut(2103)

How many seven-digit phone numbers can be formed if the first digit cannot be 0 and
a.) repetition of digits is not permitted?
9*9*8*7*6*5*4=9*(9!/3!)
b.) if repetition of digits is permitted?
9*10*10*10*10*10*10=9*10^6